# Measurement of the mean excitation energy of liquid argon

### v1

@@ Abstract
The mean excitation energy (I-value) of liquid argon is a critical input for energy estimation in neutrino oscillation experiments. It is measured to be $(205\pm 4)$ eV using the range of 402.2 MeV protons from the Fermilab Linac. This compares to the author’s recent evaluation of $(197\pm 7)$ eV based on a combination of an oscillator strength distribution analysis, gaseous argon range measurements, sparse stopping power data on solid argon, and an extrapolation of data on the effect of phase from other substances. Using all sources of information, we recommend a value of $(203.0\pm 3.2)$ eV for liquid argon, which is significantly higher than 188 eV, from ICRU-37’s gaseous argon evaluation, commonly used in Monte Carlo codes such as Geant4.
@@ 1 Introduction
Neutrino detectors such as DUNE [1] depend on knowledge of the mean excitation energy, also known as the I-value, of liquid argon to estimate neutrino energy. The I-value controls the stopping power of a substance for charged particles in the energy range where the Bethe formula is valid, $0.1\lesssim\beta\gamma\lesssim 1000$ [2]. In a liquid argon neutrino detector, the primary method to determine muon energy is by range, and the primary method of calibrating the calorimetric response is to use the minimum ionizing portion of muon tracks. Each of these relies on correct simulation that depends on liquid argon’s I-value. Typically there are no calibration points in the GeV range that can be used to mitigate an incorrect I-value. While near detectors cancel out neutrino flux and neutrino cross section uncertainties in long-baseline analyses, they do not cancel out energy reconstruction uncertainties.
@@ 2 Experimental setup
The transmittance of the materials in the beam was measured by two ceramic scintillators 0.5 m downstream of the vessel, one directly on the beam axis and the other 14 cm off-axis. Made by Advatech, the scintillators are 50 mm by 50 mm by 1 mm. They are 99.5% aluminum(III) oxide doped with chromium(III) oxide, known as chromox, with density 3.85 g/cm3 [8]. Both scintillators were observed by a single commodity video camera, a Logitech C920, positioned 1.2 m off the beam axis. The scintillators were angled 30 towards the camera with respect to the beam axis. Immediately downstream of the vessel is a stack of concrete shielding blocks with a “cave” 1 m wide by 1 m tall along the beam path that leads to the beam dump. The scintillators were positioned inside this cave, which, compared to an open configuration, slightly increases the flux of secondary particles they observe along with the primary protons.
@@ 3 Data collection
Data was obtained for 9 thicknesses of copper across a total of 31 pulses, with between 1 and 6 pulses per thickness. Unforeseen circumstances prevented further data taking. A more even distribution of data across all 20 thicknesses of copper would have been desirable, but this data is sufficient to make a measurement.
@@
Differences in the measured signal from pulse to pulse can be caused by variations in the beam intensity, the number of particles striking the camera in the pixels observing the scintillator, and effects related to the camera readout. Collectively these will be referred to as statistical uncertainty. The statistical uncertainty has been estimated via the spread of measurements using the same copper absorber thicknesses. It is strongly subdominant to the systematic errors that will be discussed in the following sections.
@@ 4 Fit procedure
The experiment was simulated with the I-value of liquid argon set to 181 eV, 189 eV, 197 eV, 205 eV, and 213 eV and with each of the 20 copper absorber thicknesses, with $10^{7}$ to $10^{8}$ protons simulated in each case, with higher statistics used for points providing more information to the fit, such that the uncertainty from Monte Carlo statistics is negligible. The normalization of the signal is a free parameter in the fit; i.e. we do not attempt any a priori calibration of the scintillator yield or camera’s efficiency. As a very good approximation, the transmission as a function of copper absorber thickness differs from one I-value to another only by a horizontal translation and a normalization (see figure 4). Therefore, the shape of this curve is used as a template in the fit which extracts the experimental I-value from the data. In the fit, there are three free parameters — the overall normalization, the left-right shift (absorber thickness), and the up-down shift (background normalization). For the off-axis fit, the background normalization is fixed to the Monte Carlo prediction, while on the on-axis fit, the data is allowed to constrain it. The results are shown in figure 5.
@@
For the on-axis result, it was found that the data prefers a sharper Bragg peak than the prediction. The largest discrepancy arises from the data taken at 8.8 mm, although the data at other positions also tends to prefer a sharper peak. We investigated whether the 8.8 mm data could be flawed, but did not find any evidence of this. For instance, it was not taken first or last, and the data from pulses immediately before and after it (11.9 mm and 15.1 mm, respectively) lies directly on the best fit line. Beam monitoring did not show the 8.8 mm pulses to be abnormal, and we could independently confirm this from the magnitude of the beam background observed in the obtained images. This discrepancy will be discussed further in sections 5.1, 5.2 and 5.3.
Given this unexpected feature, in determining the central value of our result, we exclude the data from 8.8 mm to lessen the discrepancy between data and simulation. Figure 5 shows the difference in the fit with and without this exclusion. Note that the curves almost exactly coincide for the falling edge around 16 mm. It is primarily this edge which sets the I-value, making the details elsewhere relatively unimportant to the result. The extracted I-value differs by only 0.5 eV between these two fits, and another 0.5 eV if the next group of data points at 11.1 mm is also excluded. This problem either does not occur in the off-axis data, or is obscured by the large uncertainties.
The independent results of $(205\pm 2\mathrm{(stat)})$ eV and $(212\pm 7\mathrm{(stat)})$ eV from the on-axis and off-axis scintillators, respectively, are in good agreement.
@@ 5 Systematic uncertainties
5.1 Scintillator
Figure 4: Predicted signal (with no scintillator quenching effect applied) as a function of I-value, showing that only the normalization and horizontal position of the curve change with I-value, as a good approximation. Left: the raw predictions for five I-values. Right: the same predictions, scaled and translated to match the curve for 197 eV.
@@ 5.1 Scintillator
According to the manufacturer and CERN-PS-90-42-AR [10], chromox scintillator has an extinction coefficient of $(8\pm 1)$ cm-1 for its peak light production frequency of 693 nm. However, a more recent study by Cook, Tresca and Lefferts 2014 [11] measured the attenuation length using transmission of 633 nm light to be $(0.05\pm 0.01)$ cm, i.e. an extinction coefficient of $(20\pm 4)$ cm-1, and also found that including a Birks’ quenching factor gives a better fit to proton irradiation data. Their reported best fit was for a Birks’ constant of $2.33\times 10^{-2}\,\mathrm{g\,MeV^{-1}cm^{-2}}$ and attenuation length of $0.058$ cm or $0.057$ cm (both values are quoted). This quenching constant is large compared to typical values for inorganic scintillators, although not outside what has been measured before [12].
@@
However, the effect of scintillator quenching is to flatten the Bragg peak, since quenching suppresses light production from low energy protons. This is substantially at odds with our data, which prefers a sharper Bragg peak than Monte Carlo predicts even without quenching. A resimulation of Cook’s setup was performed to investigate this discrepancy. Using the data from their Figure 2, we find that a much smaller Birks’ constant along with a shorter attenuation length fits the data somewhat better (see figure 6). Since they viewed their scintillator from the back, both parameters result in loss of light for protons with a range shorter than the scintillator thickness, and so the two effects cannot be readily separated.
For analyzing our data, we adopt a Birks’ constant of $0.25\times 10^{-2}\,\mathrm{gMeV^{-1}cm^{-2}}$, halfway between what our data prefers, zero, and the value we fit in figure 6. Varying this between zero and $0.5\times 10^{-2}\,\mathrm{gMeV^{-1}cm^{-2}}$ gives our estimate of the systematic uncertainty stemming from quenching, ±$\pm 0.8$ eV.
@@
Several rather different proposals for the value of the attenuation length are now in play. At one extreme is the manufacturer’s value of 0.125 cm, there is Cook’s value of 0.05 cm, and our value of 0.03 cm which, combined with a small value of Birks’ constant, gives a closer fit to both their data and ours. Changes in the scintillator attenuation length are almost exactly degenerate with an overall left-right shift of the predicted light curve, as it is equivalent to adding or subtracting material in the protons’ path. The difference in inferred I-value between the extremes of the possible attenuation values is 1.0 eV. We take 0.05 cm as the nominal attenuation length, assume a uniform PDF over the range and claim an uncertainty of 0.4 eV.
Figure 5: Left: Experimental data for on-axis scintillator (crosses) with a comparison to Geant4 simulations using $I=187$ eV (the ICRU-90 value for gaseous argon) and 204 eV (the best fit value). The solid red curve shows the primary fit, which excludes the data at 8.8 mm, while the dotted curve is a fit including these points. Each cluster of data points is for a single copper absorber thickness and have been given slight horizontal separations to display the uncertainties. Right: the same for the off-axis scintillator with its fit result of 212 eV, showing only the primary fit.
@@ 5.2 Beam
The uncertainty in the central energy of the beam of 0.2 MeV leads to an uncertainty on the I-value of 1.5 eV, making it one of the dominant uncertainties. In contrast, the measured I-value is quite insensitive to the details of the energy spread, with a shift of less than 0.1 eV resulting from the $(0.5\pm 0.05)$ MeV uncertainty on the beam energy width. Since our data favors a sharper Bragg peak than the Monte Carlo predicts, one could wonder if the beam energy profile is narrower than what was modeled. However, nearly all of the spread of the peak is caused by energy straggling in the experimental materials and not the initial energy profile of the beam. Even if the beam were monoenergetic, only a small fraction of the discrepancy would be explained by it.
@@
5.3 Monte Carlo code consistency
The other Monte Carlo codes predict slightly shorter or longer proton ranges than Geant4 given the same I-value, meaning that analyzing the data using them results in, respectively, higher or lower I-values. FLUKA and mars each give a result 1.6 eV higher than Geant4, while the answer from PHITS was 1.0 eV lower. As is visible from the plot, the largest difference between the codes is on the left side, before the peak, although it is the falling edge on the right side which gives the range. The PHITS result is in somewhat better agreement with the low data points at 8.8 mm, but the difference is not enough to bring data and simulation together.
Table 1: Systematic uncertainties. Materials accounting includes uncertainties on the densities and thicknesses of all materials.
Uncertainty eV
Scintillator & camera ±$\pm 1.6$
Beam energy ±$\pm 1.6$
Monte Carlo code consistency ±$\pm 1.4$
Materials accounting ±$\pm 1.2$
Multiple Coulomb scattering ±$\pm 1.0$
Aluminum & copper I-values ±$\pm 0.7$
Other physics ±$\pm 0.5$
Alignment ±$\pm 0.4$
Total systematic uncertainty ±$\pm 3.3$
Statistical uncertainty ±$\pm 1.6$
Total uncertainty ±$\pm 4\phantom{.0}$
@@ 5.4.1 Aluminum vessel
Given the specification of 6061 aluminum, there is some freedom in its elemental composition. The first-order effect for the stopping power is the material’s mean $Z/A$. The allowed range is 0.12%, giving an uncertainty in liquid argon’s I-value of 0.06 eV. The density of our particular aluminum was not measured. We take 6061 aluminum density to be $(2.69\pm 0.01)$ g/cm3 [18], and further taken as uncorrelated to the uncertainty in its elemental composition, leading to an uncertainty of 0.6 eV. The interior length of the aluminum vessel has a tolerance of 0.25 mm. The resulting uncertainty in the thickness of liquid argon leads to an I-value uncertainty of 0.3 eV. The aluminum’s density dominates the uncertainty from the aluminum vessel, which is 0.8 eV in total.

### v2

@@ Abstract
The mean excitation energy (I-value) of liquid argon is a critical input for energy estimation in neutrino oscillation experiments. It is measured to be $199^{+4}_{-5}$ eV using the range of 402.2 MeV protons from the Fermilab Linac. This compares to the author’s recent evaluation of $(197\pm 7)$ eV based on a combination of an oscillator strength distribution analysis, gaseous argon range measurements, sparse stopping power data on solid argon, and an extrapolation of data on the effect of phase from other substances. Using all sources of information, we recommend a value of $(198\pm 4)$ eV for liquid argon, which is significantly higher than 188 eV, from ICRU-37’s gaseous argon evaluation, commonly used in Monte Carlo codes such as Geant4.
[…]
@@ 1 Introduction
[…]
@@ 2 Experimental setup
The transmittance of the materials in the beam was measured by two ceramic scintillators 0.5 m downstream of the vessel, one directly on the beam axis and the other 14 cm off-axis. Made by Advatech, the scintillators are 50 mm by 50 mm by 1 mm. They are 99.5% aluminum(III) oxide doped with chromium(III) oxide, known as chromox, with density 3.85 g/cm3 [8]. Both scintillators were observed by a single commodity video camera, a Logitech C920, positioned 1.2 m off the beam axis. The camera was set to manual exposure mode with a fixed 30 frames-per-second readout. The focus was fixed and set manually. Likewise, automatic white balancing and backlight compensation were disabled.
The scintillators were angled 30 towards the camera with respect to the beam axis. Immediately downstream of the vessel is a stack of concrete shielding blocks with a “cave” 1 m wide by 1 m tall along the beam path that leads to the beam dump. The scintillators were positioned inside this cave, which, compared to an open configuration, slightly increases the flux of secondary particles they observe along with the primary protons.
@@ 3 Data collection
Data was obtained for 9 thicknesses of copper across a total of 31 pulses, with between 1 and 6 pulses per thickness. See Table 1. Unforeseen circumstances prevented further data taking. A more even distribution of data across all 20 thicknesses of copper would have been desirable, but this data is sufficient to make a measurement.
Table 1: Data taking. The 31 pulses are shown, giving the local time and copper thickness used for each.
Time Copper Time Copper Time Copper
11:18 21.5 mm 11:34 21.5 mm 12:52 17.5 mm
11:24 11.9 mm 11:35 21.5 mm 12:53 17.5 mm
11:25 11.9 mm 11:36 18.3 mm 12:54 14.3 mm
11:26 11.9 mm 11:37 18.3 mm 12:55 14.3 mm
11:27 8.8 mm 12:43 15.1 mm 12:56 14.3 mm
11:28 8.8 mm 12:47 20.7 mm 12:57 14.3 mm
11:29 8.8 mm 12:48 17.5 mm 12:58 14.3 mm
11:30 15.1 mm 12:49 17.5 mm 12:59 11.1 mm
11:31 15.1 mm 12:50 17.5 mm 13:00 11.1 mm
11:32 15.1 mm 12:51 17.5 mm 13:01 11.1 mm
11:33 21.5 mm
@@
Differences in the measured signal from pulse to pulse can be caused by variations in the beam intensity, the number of particles striking the camera in the pixels observing the scintillator, and effects related to the camera readout. These effects are not otherwise accounted for, and collectively will be referred to as statistical uncertainty. The statistical uncertainty has been estimated via the spread of measurements using the same copper absorber thicknesses. It is strongly subdominant to the systematic errors that will be discussed in the following sections.
[…]
[width=0.4]pos_03-27-00-still-00975-sub.png
@@ 4 Fit procedure
The experiment was simulated with the I-value of liquid argon set to 181 eV, 189 eV, 197 eV, 205 eV, and 213 eV and with each of the 20 copper absorber thicknesses, with $10^{7}$ to $10^{8}$ protons simulated in each case, with higher statistics used for points providing more information to the fit, such that the uncertainty from Monte Carlo statistics is negligible. The normalization of the signal is a free parameter in the fit; i.e. we do not attempt any a priori calibration of the scintillator yield or camera’s efficiency. As a very good approximation, the transmission as a function of copper absorber thickness differs from one I-value to another only by a horizontal translation and a normalization (see figure 4). Therefore, the shape of this curve is used as a template in the fit which extracts the experimental I-value from the data. In the fit, the parameter of interest is the left-right shift (absorber thickness). Additional parameters are the up-down shift (background normalization), the overall normalization, and three free parameters for several systematic effects, discussed below. Because the off-axis data has large statistical uncertainties, the background normalization is fixed to the Monte Carlo prediction, and three systematics parameters are fixed to the best fit values found in the on-axis fit. The results are shown in figure 5.
[…]
Figure 4: Predicted signal (energy deposition) as a function of I-value, showing that only the normalization and horizontal position of the curve change with I-value, as a good approximation. Left: the raw predictions for five I-values. Right: the same predictions, scaled and translated to match the curve for 197 eV.
[…]
Figure 5: Left: Experimental data for on-axis scintillator (crosses) with a comparison to Geant4 simulations using $I=187$ eV (the ICRU-90 value for gaseous argon) and 199 eV (the best fit value). The solid red curve shows the best fit including systematics effects of the scintillator and camera, while the dotted curve shows the expectation for 199 eV if detector response were directly proportional to energy deposition. Each cluster of data points is for a single copper absorber thickness and have been given slight horizontal separations to display the uncertainties. Right: the same for the off-axis scintillator, showing only 187 eV and the best fit.
@@
It was found that the on-axis data prefers a sharper Bragg peak than the prediction under the naive assumption that the detector signal is simply proportional to energy deposition. This is particularly pronounced with the data taken at 8.8 mm, although the data at other positions also tends to prefer a sharper peak. We investigated whether the 8.8 mm data could be flawed, but did not find any evidence of this. For instance, it was not taken first or last (see Table 1), and the data from pulses immediately before and after it (11.9 mm and 15.1 mm, respectively) are not similarly lower than expected. Beam monitoring did not show the 8.8 mm pulses to be abnormal, and we could independently confirm this from the magnitude of the beam background observed in the obtained images.
We investigated several detector effects and determined that the observed shape could be explained via a combination of quenching in the scintillator and different emission spectrum of the scintillator as a function of particle dE/dx, convoluted with the camera’s spectral response. These effects, and systematic uncertainties associated with them, will be discussed at length in sections 5.1 and 5.2.
The statistically independent results of $(198\pm 1\mathrm{(stat)})$ eV and $(208^{+6}_{-7}\mathrm{(stat)})$ eV from the on-axis and off-axis scintillators, respectively, are in good agreement. The statistical uncertainty is defined as the shift necessary to produce a $\Delta\chi^{2}$ of 1.0 when only the mean ionization energy and the overall normalization are allowed to vary.
@@ 5 Systematic uncertainties
Table 2: Systematic uncertainties. Each category is explained in detail in the listed section.
Uncertainty eV Section
Scintillator & camera ${}^{+2.5}_{-4}$ 5.1
Beam energy ±$\pm 1.6$ 5.2
Materials accounting ±$\pm 1.2$ 5.3
Multiple Coulomb scattering ±$\pm 1.0$ 5.4
Monte Carlo code consistency ±$\pm 0.9$ 5.5
Aluminum & copper I-values ±$\pm 0.7$ 5.6
Other physics ±$\pm 0.5$ 5.7
Alignment ±$\pm 0.4$ 5.8
Total systematic uncertainty ${}^{+4}_{-5}$
Statistical uncertainty ±$\pm 1.3$
Total uncertainty ${}^{+4}_{-5}$
5.1 Scintillator and camera
@@
According to the manufacturer and CERN-PS-90-42-AR [10], chromox scintillator has an extinction coefficient of $(8\pm 1)$ cm-1 for its peak light production frequency of 693 nm. A more recent study by Cook, Tresca and Lefferts 2014 [11] measured the attenuation length using transmission of 633 nm light — substantially below the peak emission — to be $(0.05\pm 0.01)$ cm, i.e. an extinction coefficient of $(20\pm 4)$ cm-1, and also found that including a Birks’ quenching factor gives a better fit to proton irradiation data. Their best fit was for a Birks’ constant of $2.33\times 10^{-2}\,\mathrm{g\,MeV^{-1}\,cm^{-2}}$ and attenuation length of $0.058$ cm or $0.057$ cm (both values are quoted). This quenching constant is large compared to typical values for inorganic scintillators, although not outside what has been measured before [12].
@@
The effect of scintillator quenching is to flatten the Bragg peak, since quenching suppresses light production from low energy protons. This is substantially at odds with our data, which prefers a sharper Bragg peak than the simulation predicts even without quenching. A resimulation of Cook’s setup was performed to investigate this discrepancy. Using the data from their Figure 2, we find that a much smaller Birks’ constant along with a shorter attenuation length fits the data slightly better (see figure 6). Since they viewed their scintillator from the back, both parameters result in loss of light for protons with a range shorter than the scintillator thickness, and so the two effects cannot be readily separated. On the other hand, both their direct attenuation measurement and the manufacturer’s statement disfavor a shorter length.
[…]
[width=0.5122]fig-cook.pdf
We choose to take $(2.33\pm 0.50)\times 10^{-2}\,\mathrm{g\,MeV^{-1}\,cm^{-2}}$ as the nominal value and uncertainty of Birks’ constant. We estimate this uncertainty from the amount of freedom in the quenching allowed by Cook’s attenuation measurement. The uncertainty in the I-value incurred by the quenching turns out to be negligible because its effect is relatively well-constrained and, in any case, it is degenerate with the effect we will discuss next.
To explain the discrepancy between our data and the expected shape of the Bragg peak, we suggest that the scintillation spectrum of chromox shifts with the dE/dx of protons. Such a shift has been observed when comparing emission from x-ray and alpha irradiation in zinc oxide and zinc sulfide scintillators [13], in which the highly ionizing alphas result in a spectrum shifted roughly 10 nm down. Since we observed the chromox emission through both an infrared filter (“hot mirror”) and a red filter in our camera, we are sensitive to the convolution of the emission spectrum and transmission efficiency. A similar 10 nm shift would cause a 50% change in the signal (see figure 7) because of the sharp edge of the infrared filter, while a 20 nm shift would cause a 200% change.
[…]
Figure 7: Left: chromox emission spectrum from Ref. [14], C920 infrared filter transmittance from Ref. [15], the typical product of red filter transmittance and CMOS efficiency for a digital camera, and the overall product of emission and efficiency. Right: resulting overall efficiency, relative to nominal, as a function of a uniform shift of the chromox spectrum.
A second effect that may skew the signal as a function of dE/dx is a varying scintillator decay time. It has been observed that x-rays produce long afterglows in chromox [14] and also that the decay time when irradiated with 50 keV He+ is much longer (1.6 s to fall 90%) than when irradiated with 1.06 MeV protons (72 ms) or 1.65 MeV He+ (60 ms) [16], all of which are much larger than the manufacturer’s figure of 3.4 ms, which perhaps refers to the decay time for high energy proton irradiation.
Since we lack specific data for either the spectral shift or the scintillator decay time as a function of dE/dx, we make a generic assumption that there is one efficiency for low-dE/dx particles and another for high-dE/dx particles, with some transition region between these. We allow the fit to determine the cutoff dE/dx, the width of the transition, and the ratio of efficiencies. We limit the ratio of efficiencies to between 1:1 and 1:3, corresponding to a spectral shift in the chromox of between zero and 20 nm. The data has relatively little power to simultaneously constrain all three parameters, leading to a large systematic uncertainty. The fit prefers an abrupt transition at 60 MeV/cm with an efficiency ratio of 1:2.3, but given the degeneracies between parameters, these specifics should not be taken too literally.
The effect of higher efficiency for higher dE/dx is to increase the signal on the falling edge of the Bragg peak while lowering it on the leading edge. This is illustrated by the difference between dotted and solid lines in figure 5, left. The change allows both the data in the peak and that at 8.8 mm to be well-fit. Unfortunately, it is somewhat correlated to an overall right-left shift and so incurs a large systematic uncertainty of ${}^{+3.2}_{-4}$ eV.
Finally, we take a systematic uncertainty for the attenuation length itself. The attenuation lengths from the manufacturer of 0.125 cm for 693 nm and from Cook of 0.05 cm for 633 nm may both be correct, since similar materials are known to have rapidly changing attenuation lengths in this wavelength region [17], with shorter lengths at shorter wavelengths. As can be seen from figure 7, the visible signal for us is split roughly evenly between the sharp peak at 693 nm and the broad peak at 620–680 nm. The relative contributions are not well known because of the possible wavelength-shifting effect discussed above, nor do we know how the attenuation length evolves between the two measured wavelengths.
@@
Changes in the scintillator attenuation length are almost exactly degenerate with an overall left-right shift of the predicted light curve, as it is equivalent to adding or subtracting material in the protons’ path. We consider the extreme cases of all of the light having an attenuation length of 0.05 cm and all of it having an attenuation length of 0.125 cm. The difference in inferred I-value between these extremes is 0.7 eV. We take the midpoint as the nominal result, assume a uniform PDF over the range and claim an uncertainty of 0.2 eV.
@@ 5.2 Beam
The uncertainty in the central energy of the beam of 0.2 MeV leads to an uncertainty on the I-value of 1.5 eV. In contrast, the measured I-value is quite insensitive to the details of the energy spread, with a shift of less than 0.1 eV resulting from the $(0.5\pm 0.05)$ MeV uncertainty on the beam energy width. It is notable that nearly all of the spread of the Bragg peak is caused by range straggling in the experimental materials and not the initial energy profile of the beam.
@@ 5.3.1 Aluminum vessel
Given the specification of 6061 aluminum, there is some freedom in its elemental composition. The first-order effect for the stopping power is the material’s mean $Z/A$. The allowed range is 0.12%, giving an uncertainty in liquid argon’s I-value of 0.06 eV. The density of our particular aluminum was not measured. We take 6061 aluminum density to be $(2.69\pm 0.01)$ g/cm3 [19] at room temperature — $(2.72\pm 0.01)$ g/cm3 at 87 K — and further assume the density uncertainty is uncorrelated to the uncertainty in its elemental composition, leading to a total uncertainty of 0.6 eV. The interior length of the aluminum vessel has a tolerance of 0.25 mm. The resulting uncertainty in the thickness of liquid argon leads to an I-value uncertainty of 0.3 eV. The aluminum’s density dominates the uncertainty from the aluminum vessel, which is 0.8 eV in total.